The printtim package^{∗}

Martin Schröder

Friedrich-Humbert-Straße 124 D-28759 Bremen MS@Dream.HB.North.DE (INTERNET)

1995/01/24

Abstract This package defines a new command \PrintTime which prints out the

time the document was processed.

Contents 1 Introduction

This package provides the command \PrintTime which prints out the time the document was processed.

2 Usage

\PrintTime prints out the current time: “20:03”.

3 Options

The package has no options.

4 Required packages

The package does not require any further packages.

∗The version umber of this file is v1.00, last revised 1995/01/24.

The name printtim is a tribute to the 8+3 file-naming convention of certain “operating systems”; strictly speaking it should be printtime.